Questions

You must show your work to obtain full credit.

1- Which weighs more,8.0 mL of Pb ( density = 11.3g/mL) or 41.0 g of Hg?

2- Calculate the mass percent of N in N2O5 (MM of N2O5 = 108.01 g/mol):

Mass percent N =

3- The empirical formula for iso-octane,C8H18 is

4- (SF) what is the density of Zirfinium if you know that 1.6 mL weighs 1.40g?

5-(SF) How many atoms are present in 354.53 amu of silicon?

6-How many molecules are present in 354,53 amu of Cl2.?

7- (units) Calculate the molar mass of di-silane, Si2H6.?

8- Follow the steps in 8-12 to calculate the empirical formula of compound of Sulfur and Oxygen when you this compound contains 50.05% S by mass

Mass Of O in a 100.00g sample =

Mols of O =

Mols of S=
Answer #1

1) Mass of 8.0mL Pb = Volume x density = 8.0mL x 11.3g/mL=90.4 g

Sp Pb is heavier than 41 g of Hg

2)N2O5 molar mass =108.01 g/mol

mass due to Nitrogen = 28 g

thus mass % of N = mass of N x100 /molar mass

= 28x100/108.01 = 25.92 %

3)empirical formula is the simplest ratio of atoms in a molecule.

The M.F. of isooctane is C8H18 and the simplest ratio = empirical formula = C8H18

4) density = mass / volume

= 1.40g/1.6 mL

= 0.875 g/mL

5) Si atomic mass = 28.085amu.

thus 28amu has one Si atom, then 354.53 amu has  

Si atoms = 354.53/28.085=12.62 atoms